A person standing at the junction of two straight paths represented by the equations 2x-3y+4=0 and 3x+4y-5=0wants to reach the path whose equation is 6x-7y+8=0 in least time. Find equation of path that he should follow.

Grade:12th Pass

1 Answers

akash dutta
5 years ago
2x-3y+4=0
3x+4y-5=0
 
By solving both of these
y=22/7
x=-1/17
 
Slope of straight line=6/7
 
So the eqn of the line straight to it = ???
Its slope is -7/6
 
So by putting the values in the formula 
The required eqn of the line is
102x+102y-125=0